OpenStudy (anonymous):

In triangle STU, s = 9 cm, t = 15 cm, and m∠U = 37º. Find u.

OpenStudy (anonymous):

the answer choice is 7.4 cm 8.6 cm 9.5 cm this was my guess? is it right? 10.3 cm

OpenStudy (anonymous):



OpenStudy (anonymous):

sinS/9=sin(143-S)/15=sin37/u 15sinS=9sin(143-S) 15sinS=9(sin143*cosS-cos143*sinS 15sinS=9(0.602*cosS+0.799sinS) 7.809sinS=5.418cosS tanS=0.694 =>S=34.76

OpenStudy (anonymous):

did you get this answer right?